OpenStudy (princeevee):

solve for c -1/4c = -20 5 -80 -5 80

pooja195 (pooja195):

Multiply by the reciporcal So just flip the fraction -1/4=. -4/1 So... -20 x -4 =?

OpenStudy (anonymous):

C=80

OpenStudy (princeevee):

but i got - 80

pooja195 (pooja195):

A negative times a negative is a positive.

OpenStudy (anonymous):

Yours wrong

OpenStudy (princeevee):

oh ok

OpenStudy (princeevee):

2 more?

OpenStudy (anonymous):

Yeah!

pooja195 (pooja195):

--20 --4 Both numbers are negative when you multiply 2 negative numbers the answer will be positive so its 80 NOT -80

OpenStudy (anonymous):

Give us some more to do but don't show us de answer lol

OpenStudy (princeevee):

3/4x=9 1/4 4/3 4 -3/4

OpenStudy (anonymous):

X=12

pooja195 (pooja195):

Multiply by the reciporcal So just flip the fraction 3/4=4/3 So... 9 x 4/3 = ?

OpenStudy (princeevee):

well i am not good with fractions